OpenStudy (praxer):

What is mutually exclusive events, independent events and The multiplication theory of probability ?

OpenStudy (praxer):

@kc_kennylau please help....

OpenStudy (anonymous):

to events \(A\) and \(B\) are "mutually exclusive" if \(A\cap B=\emptyset\)

OpenStudy (anonymous):

it means they cannot happen at the same time like if you roll two dice the total can't be even and 7

OpenStudy (anonymous):

two events \(A\) and \(B\) are "independent" means \[P(A|B)=P(A)\]

OpenStudy (isaiah.feynman):

Two events are mutually exclusive if they can't have the same outcome. Two events are independent when the probability of one doesn't affect the outcome of the other.

OpenStudy (isaiah.feynman):



OpenStudy (anonymous):

in other words, knowing that \(B\) has occurred gives no additional information about the probability of \(A\) for example, if you flip a coin twice and know the first toss is heads, it does not change the probability of the second toss being heads

OpenStudy (anonymous):

"multiplication" will be \[P(A\cap B)=P(A)\times P(B|A)\]

OpenStudy (anonymous):

if \(A\) and \(B\) are independent, this means \[P(A\cap B)=P(A)\times P(B)\]

OpenStudy (anonymous):

since if \(A\) and \(B\) are independent, then \(P(B|A)=P(B)\)

OpenStudy (praxer):

Since in mutually exclusive events both cannot occur at the same time we add it right ? In independent events since both are independent or one doesn't affect the other we multiply it... Am i correct ?